OpenStudy (anonymous):

Help please I keep getting a negative number. Solve the system of equations algebraically using (substitution or elimination) y=40x+95 y=50x+30

OpenStudy (phi):

what did you get ?

OpenStudy (anonymous):

What are the options? because their are multiple answers to this. @HappyBunny

OpenStudy (anonymous):

It was a long decimal it didn't make sense so i erased it

OpenStudy (phi):

do you want to use substitution or elimination ?

OpenStudy (anonymous):

elimination

OpenStudy (phi):

in that case subtract the two equations y=40x+95 - ( y=50x+30) -------------- y-y = 40x - 50x +95 -30 can you simplify that ?

OpenStudy (anonymous):

-10x+65 ?

OpenStudy (phi):

ok, but you should write the whole equation. on the left side you have y-y which is 0 0 = -10x +65 now add +10x to both sides

OpenStudy (anonymous):

ohhh! x=75 and then I plug it into y=45x+95 correct?

OpenStudy (phi):

? 0 = -10x +65 +10x = +10x - 10x + 65 simplify

OpenStudy (phi):

on the right you simplify 10x - 10x to 0

OpenStudy (anonymous):

so then it would just be 10x=65?

OpenStudy (phi):

yes. next divide both sides by 10 \[ \frac{10x}{10} = \frac{65}{10} \]

OpenStudy (phi):

you know 10/10 is 1

OpenStudy (anonymous):

x=13/2 ?

OpenStudy (phi):

yes, or x=6.5

OpenStudy (anonymous):

Then I just plus the x into one of the equations to find Y Right?

OpenStudy (phi):

yes. what do you get ?

OpenStudy (anonymous):

I did y= 40(6.5)+95 and 260+95 y= 355

OpenStudy (phi):

yes. so the answers are x= 13/2 and y=355 you can use the other equation to double check the answer

OpenStudy (anonymous):

Thank you I've been stuck on this question for a long time

OpenStudy (phi):

yw